OpenStudy (anonymous):

the expression √ 72-3√ 2 written in simplest radical form is..?

OpenStudy (anonymous):

you have to break down whats under the radical so they match the other radical and pull the square out then subtract it out

OpenStudy (anonymous):

did that make sense?

OpenStudy (anonymous):

not really lol

OpenStudy (anonymous):

simplyy 72 by thing of it as 36*2 sqroot of 36 is 6 so sqroot of 72 is 3 sqroot2 6 sqrt2 - 3sqrt2 = 3sqrt2

OpenStudy (anonymous):

\[\sqrt{72}=\sqrt{(36)(2)}=6\sqrt{2}\]

OpenStudy (anonymous):

sorry sqrt72 is 6sqrt2

OpenStudy (anonymous):

ohhhhhhhhhhhhhh i get it haha thankyou:)

OpenStudy (anonymous):

no problem best of luck to you